A woman is q years old while her son is p years old. The sum of their ages is equal to twice the difference of there ages, the product of their ages is 675. Find the ages of woman and her son?
Answer by htmentor(1343)

Sum of their ages is twice the difference:
p + q = 2(q - p) -> p + q = 2q - 2p -> q = 3p
Product of their ages is 675:
p*q = 675
Substitute the value for q above:
p*3p = 675
3p^2 = 675
p = 15
So the son is 15 and the mother is 45 yrs old